A 1200-kg car is being raised with a constant acceleration of 2.53 m/s2 by a crane, using a 20-m long steel cable that is 1.5 cm in diameter. Young's modulus for steel is 2.0 × 1011 N/m2. What is the change in length of the cable caused by lifting the car?
Answer #1

Tension in wire T-mg=ma m T= mig+a)=1200 (9.8t 2053) = 14796 N Ya stress strain TA ALL LL = TL = 14796420 AY TIX 2 x 2x1011 )